OpenStudy (anonymous):

Find an equation for the tangent line at the point where x=2 on the graph of the function f(x)=5^x/2

OpenStudy (anonymous):

The slope of the tangent line is the derivative evaluated at the point. Do you know how to find the derivative of your function? (And is that \(\frac{5^x}{2}\) or \(5^{\frac{x}{2}}\)?

OpenStudy (anonymous):

no, I couldn't figure out how to find the derivative. Btw the function is 5^(x/2)

OpenStudy (anonymous):

The derivative of \(a^x\) is \(a^x\ln{a}\).

OpenStudy (anonymous):

So the derivative of your function is \(\frac{1}{2}5^{\frac{x}{2}}\ln{5}\)

OpenStudy (anonymous):

where did the 1/2 come from?

OpenStudy (anonymous):

The derivative of \(\frac{x}{2}\) is \(\frac{1}{2}\).

OpenStudy (anonymous):

(You need to use the chain rule in the exponent.)

OpenStudy (anonymous):

ok thank you!

OpenStudy (anonymous):

how do you find the equation?

OpenStudy (anonymous):

Find f(2), you'll get the coordinates of a point. Find f'(2), you'll get the slope of the tangent line. Equation of the tangent line can thus be obtained by the point-slope form.